40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 25 16:05:51 2022, Yehonathan, Update, SUS, Trying to fix some SUS PR2_Step_Response_Test_2022-08-25_16-23.pdfPR3_Step_Response_Test_2022-08-25_18-37.pdf
    Reply  Tue Sep 6 17:39:40 2022, Paco, Update, SUS, LO1 LO2 AS1 AS4 damping loop step responses LO1_Step_Response_Test_2022-09-06_17-19.pdfLO2_Step_Response_Test_2022-09-06_17-30.pdfAS1_Step_Response_Test_2022-09-06_17-53.pdfAS4_Step_Response_Test_2022-09-06_18-16.pdf
       Reply  Tue Nov 8 10:23:42 2022, Anchal, Update, SUS, MC3 damping loop step responses MC3_Step_Response_Test_2022-11-08_09-48.pdfMC3_Step_Response_Test_2022-11-08_10-19.pdf
Message ID: 17105     Entry time: Thu Aug 25 16:05:51 2022     Reply to this: 17133
Author: Yehonathan 
Type: Update 
Category: SUS 
Subject: Trying to fix some SUS 

I tried to lock the Y/X arms to take some noise budget. However, we noticed that TRX/Y were oscillating coherently together (by tens of percent), meaning some input optics, essentially PR2/3 are swinging. There was no way I could do noise budgeting in this situation.

I set out to debug these optics. First, I notice side motion of PR2 is very weakly damped .

The gain of the side damping loop (C1:SUS-PR2_SUSSIDE_GAIN) was increased from 10 to 150 which seem to have fixed the issue. Attachment 1 shows the current step response of  the PR2 DOFs. The residual Qs look good but there is still some cross-couplings, especially when kicking POS. Need to do some balancing there.

PR3 fixing was less successful in the beginning. I increased the following gains:

C1:SUS-PR3_SUSPOS_GAIN: 0.5 -> 30

C1:SUS-PR3_SUSPIT_GAIN: 3 -> 30

C1:SUS-PR3_SUSYAW_GAIN: 1 -> 30

C1:SUS-PR3_SUSSIDE_GAIN: 10 -> 50

But the residual Q was still > 10. Then I checked the input matrix and noticed that UL->PIT is -0.18 while UR->PIT is 0.39. I changed UL->PIT (C1:SUS-PR3_INMATRIX_2_1) to +0.18. Now the Q became 7. I continue optimizing the gains.

Was able to increase C1:SUS-PR3_SUSSIDE_GAIN: 50 -> 100.

Attachment 2 shows the step response of PR3. The change of the entry of the input matrix was very ad-hoc, it would probably be good to run a systematic tuning. I have to leave now, but the IFO is in a very misaligned state. PR3/2 should be moved to bring it back.

Attachment 1: PR2_Step_Response_Test_2022-08-25_16-23.pdf  47 kB  Uploaded Thu Aug 25 17:51:52 2022  | Hide | Hide all
PR2_Step_Response_Test_2022-08-25_16-23.pdf PR2_Step_Response_Test_2022-08-25_16-23.pdf PR2_Step_Response_Test_2022-08-25_16-23.pdf PR2_Step_Response_Test_2022-08-25_16-23.pdf
Attachment 2: PR3_Step_Response_Test_2022-08-25_18-37.pdf  47 kB  Uploaded Thu Aug 25 19:40:05 2022  | Hide | Hide all
PR3_Step_Response_Test_2022-08-25_18-37.pdf PR3_Step_Response_Test_2022-08-25_18-37.pdf PR3_Step_Response_Test_2022-08-25_18-37.pdf PR3_Step_Response_Test_2022-08-25_18-37.pdf
ELOG V3.1.3-